OpenStudy (anonymous):

a missile is launched from the ground. its height, h(x), can be represented by a quadratic function in terms of time, x, in seconds. after 1 second, the missile is 103 feet in the air; after 2 seconds, it is 192 feet in the air. find the height, in feet, of the missile after 5 seconds in the air. help!

OpenStudy (anonymous):

Supposing the missile is at height 0 at time 0. Then you have three points: (0, 0) (1, 103) and (2, 192) These points fit the quadratic function h(x) = 110x - 7x^2 So at 5 seconds h(5) = 110(5) - 7(5^2) = 375 feet

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

@vegas14 how could u trace out that form???????? trail and error?????